Circular No. 7578

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 C1 L. Manguso, Lincoln Laboratory, reports the discovery of an apparent

comet by LINEAR. Confirming CCD observations by G. Hug (Eskridge, KS) reveal a condensed coma and a suggestion of a faint, broad tail in p.a. ∼ 325^{o}. Available astrometry:

2001 UT α_{2000} δ_{2000} m_{1} Observer
Feb. 1.48039 14^{h}41^{m}46.^{s}55 −10^{o}15′14′′.8 18.9 LINEAR
1.49581 14 41 46.66 −10 15 21.8 19.5  _{”}
1.52622 14 41 46.75 −10 15 32.2 18.6  _{”}
2.49528 14 41 49.06 −10 19 51.8 18.3 Hug
2.50030 14 41 49.08 −10 19 52.8 17.9  _{”}
2.51013 14 41 49.11 −10 19 55.0 18.3  _{”}

COMET C/1999 T1 (McNAUGHT-HARTLEY) M. J. Mumma, N. Dello Russo, and M. A. DiSanti, Goddard Space

Flight Center, NASA; K. Magee-Sauer, Rowan University; R. Novak, Iona College; and A. Conrad and F. Chaffee, W. M. Keck Observatory, report: “Water and CO were detected simultaneously near 4.67 μm on Jan. 13.7 UT in observations made at the NASA Infrared Telescope Facility (+ CSHELL). Three lines in the 1-0 band of CO (R0, R1, and P2) and two lines in the ν_{3}-ν_{2} band of H_{2}O yielded production rates (× 10^{27} molecules s^{−1}) of 14 for CO and 82 for H_{2}O. On Jan. 14.7, C_{2}H_{6} (ν_{7}, nine Q-branches), CH_{3}OH (ν_{3} Q-branch and other lines), and OH ‘prompt’ emission were detected using NIRSPEC at the W. M. Keck Observatory. The waterproduction rate derived from OH ‘prompt’ emission (P12.5 1^{−} and 1^{+}, near 3042 cm^{−1}) was 160 (calculated using g-factors from comet C/1999 H1), and other production rates were 2.7 for CH_{3}OH and 1.1 for C_{2}H_{6}. A residual 2σ signal was seen at the expected position of CH_{4} R0 (ν_{3} band), and this is formally consistent with a 3σ upper limit of 2.5 for CH_{4}. The mixing ratios are then H_{2}O:CO = 100:17 on Jan. 13.7, and H_{2}O:CH_{3}OH:C_{2}H_{6}:CH_{4} = 100:1.7:0.65:(<1.6) on Jan. 14.7. A rotational temperature of 70 K was adopted for all species on both dates. Multiple lines of HCN and C_{2}H_{2} were detected on Jan. 14, and quantitative analysis is in progress. The CO mixing ratio in this comet is similar to that found for native CO in comets C/1996 B2 and C/1995 O1, but it is much higher than those found for comets C/1999 H1 and C/1999 S4. Abundances of C_{2}H_{6} and CH_{3}OH are similar to those in comets C/1996 B2, C/1995 O1, and C/1999 H1.”

2001 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green